# Red Teaming AI Red Teaming

### v1

@@ 3.1.2 Design
Red teams should scrutinize the architecture proposed for an AI system for potential failure modes. How might different components interact in unexpected ways? What happens when the system encounters edge cases or operates outside its intended parameters? How could data flows be corrupted, manipulated, or intercepted? What are the implications of choosing particular algorithms, frameworks, or third-party components? [19]
Human-AI interaction design deserves particular attention. Red teams should challenge assumptions about user behavior, exploring how people might game the system, develop overreliance, or use it in ways that compound rather than mitigate risks [2]. They should also examine how the system communicates uncertainty, handles errors, and maintains user agency.
@@ 3.1.3 Data
Red teams should challenge assumptions about data representativeness and quality. Training datasets frequently underrepresent certain populations or edge case scenarios, creating systematic vulnerabilities. Historical biases embedded in the data perpetuate and amplify existing inequities, particularly when combined with algorithmic optimization [29]. Data distribution shifts over time can adversely impact system performance in ways that may not be immediately apparent.
@@ 3.1.4 Development
Red teaming during the development of an AI system focuses on how implementation choices create systemic vulnerabilities and how development practices themselves introduce risks. Red teams should examine the security of the development environment, with particular attention to supply chain risks. Code and model artifacts require robust security and versioning practices to prevent unauthorized modification. Dependency management represents a significant attack vector, as compromised third-party libraries can introduce vulnerabilities throughout the system [33]. Supply chain attacks targeting development tools can compromise entire development pipelines.
@@ 3.1.5 Deployment
User interaction patterns in production often differ significantly from development assumptions, creating new attack vectors and safety risks [21]. Real users exhibit behaviors that can expose edge cases leading to incorrect outputs or system failures. The scale and diversity of production usage can reveal reliability issues that only manifest under specific load patterns, while safety-critical applications must account for (benign and adversarial) user behaviors that could lead to harmful outcomes.
@@ 3.1.6 Maintenance
Model drift is a fundamental threat to system reliability and safety beyond security concerns [10]. AI system performance typically degrades over time as real-world conditions diverge from training assumptions, potentially leading to incorrect decisions in safety-critical contexts. This degradation can manifest as gradual accuracy loss, increased (statistical) bias and variance, or failures in edge cases that compromise reliability, security, and safety. The decision of when to retrain models also involves balancing multiple risks: maintaining degraded performance, introducing new failure modes through updates, and ensuring continuity of service.
@@ 3.2 Micro-level (Model) Red Teaming
Inie et al. [20] proposed a grounded theory of LLM red teaming, based on interviews with a diverse group of AI and security practitioners, in a recent paper. We follow them to summarize best practices for red teaming the generative models underlying AI systems.
@@
To achieve these goals, Inie et al. [20] developed a taxonomy of 12 strategies and 35 techniques, organized into five categories that combine technical knowledge with creative problem solving. Conceptually, this structure is reminiscent of the Tactics, Techniques, and Procedures (TTP) framework in cybersecurity.
@@
Effective model red teaming requires incorporating diverse user perspectives and community voices into the evaluation process [45]. Red teaming should go beyond simply testing at scale using adversarial prompt variations to consider how different communities experience AI systems and what constitutes harm from their perspectives. This approach recognizes that AI vulnerabilities often manifest differently across demographic groups and use contexts, making diverse participation essential for comprehensive risk discovery.
@@
Macro-level and micro-level red teaming are complementary approaches that together provide comprehensive coverage of AI system risks. Macro-level insights inform micro-level testing priorities, while micro-level findings inform macro-level risk assessments—technical vulnerabilities have implications for system architecture and operational procedures. Effective AI red teaming requires both perspectives applied consistently throughout the development lifecycle [50]. Ideally, it should be a collaboration that brings together diverse expertise, both technical and non-technical, to challenge assumptions and identify blind spots in stages of the lifecycle [53].
@@ 4 A Systems Perspective
Certain aspects of AI red teaming go beyond the implementation guidelines of mostly technical components that we presented in the last section. There are broad principles that guide the effective coordination of macro-level system red teaming and micro-level model red teaming, and their integration into the broader socio-technical environment the AI system interacts with. In this section, we draw upon established systems theory concepts, recent advances in understanding agentic AI systems, and coordinated disclosure frameworks to offer strategic recommendations for orchestrating effective AI red teaming. We call this Meta-level Red Teaming, aimed at addressing novel and emergent risks arising from complex interactions between technical components, organizational contexts, and deployment environments.
@@ 4.1 Recommendation 1: Adopt Systems-Theoretic Perspectives
A recent position paper by Miehling et al. [30] argues that current agentic AI development is overly focused on individual model capabilities, often ignoring broader emergent behavior, leading to a significant underestimation in true capabilities and unaccounted for risks. This insight directly applies to red teaming: In addition to testing components in isolation, organizations should implement system-level critical thinking that examines how undesired functional agency may emerge from component interactions inside an agentic AI system.
Agentic systems, consisting of agents that iteratively interact with humans and other agents to achieve specified tasks, possess properties that are amenable to a system-level analysis. At the most granular level, an agent contains an internal reason-act-sense-adapt loop. This loop feeds and is fed by feedback loops at higher levels, namely at the agent-human interface, the agent-agent interface, and the agent-environment interface [30]. Agentic AI red teaming must map these interaction patterns to identify vulnerabilities that emerge only from multiagent coordination, human-AI collaboration, and environmental adaptation.
@@ 4.2 Recommendation 2: Pair Red Teaming with Testing
To ensure comprehensive assessment of AI systems throughout their lifecycle, AI red teaming must be positioned within the broader context of Test, Evaluation, Verification, and Validation (TEVV). As CISA notes, AI red teaming is fundamentally a subset of AI TEVV, which itself must be grounded in established software TEVV practices while accounting for AI-specific characteristics [9].
Conventional AI benchmarking and evaluation practices exhibit systemic limitations for exploring, navigating, and resolving the human and societal factors that occur in real world deployment [42]. Keeping this in mind, AI red teaming should address both first-order effects (immediate system output, such as accuracy, toxicity, or bias) and second-order effects (long-term outcomes and consequences of AI use, such as shifts in user behavior, societal ramifications, and workforce transformations).
This requires expanding red teaming beyond static, single-turn automated testing to include paradigms that capture realistic interaction patterns between diverse groups of users and AI technology in context [42]. Red teams should design evaluation scenarios that account for contextual deployment factors and emergent behaviors that surface only through sustained real-world interaction.
@@ 4.3 Recommendation 3: Coordinated Disclosure Infrastructure
Similar to software vulnerabilties, AI failures may be transferable between models and systems, and multiple stakeholders can play a role in their mitigation [25]. For proper management of this situation, organizations should implement coordinated disclosure mechanisms with standardized reporting processes. Similarly to cybersecurity, AI disclosure coordination centers can route flaw reports across the AI ecosystem through tagged subscriptions, for instance, Meta subscribing to “Llama 3.3” tags or government agencies subscribing to “Impacts: Cybersecurity” tags [25]. This process should include standardized flaw report templates, rules of engagement for evaluators, and extending safe harbor protection to AI red teamers and offensive researchers following good faith protocols.
@@ 4.4 Recommendation 4: Design Bidirectional Feedback Loops
Building on systems theory principles, organizations should establish formal processes for macro-micro information exchange throughout the red teaming lifecycle [24]. Macro-level insights about deployment environments and organizational contexts should systematically inform micro-level testing priorities, while micro-level vulnerability discoveries should trigger macro-level risk reassessments.
For agentic AI systems specifically, this means testing how micro-level model behaviors interact with macro-level tools, deployment contexts, and multi-agent interactions. Enhanced cognition arises due to the agent’s interaction with the environment via tools, effectively acting as the “sensorimotor” interface that enables the agent to perceive and manipulate its environment [30].
@@ 4.5 Recommendation 5: Threat Modeling for Emergent Risks
Traditional threat modeling focuses on individual system components, but AI systems require multilevel threat modeling that maps attack vectors across technical, social, and emergent behavioral dimensions. Systems theory stresses how each component of a system must be understood in terms of both its individual definition and its contribution to the larger system behavior [30].
For multi-agent systems specifically, threat modeling must account for the complex interactions between multiple AI agents, their coordination mechanisms, and emergent behaviors that arise from agent collaboration [34]. AI red teams—and TEVV frameworks—should develop threat models that explicitly represent how adversaries might exploit emergent properties that arise from system interactions. For example, allowing agents to form predictions (with associated confidences) of concepts in their environment, and additionally facilitating communication of these uncertainties to other agents, can allow the formation of shared representations and the emergence of metacognitive awareness [30]. Such emergent metacognitive capabilities could be exploited in ways that are not apparent from the testing of individual components.
@@ 4.6 Recommendation 6: Monitor for Behavioral Drifts
Systems theory emphasizes that complex systems exhibit dynamic behaviors that change over time [44]. AI red teaming should be an ongoing process that looks out for novel flaws and vulnerabilities arising from system evolution and environmental changes.
For agentic AI systems, this is particularly critical because mechanisms for enhanced agent cognition, emergent causal reasoning ability, and metacognitive awareness can develop through environmental interaction [30]. Organizations should implement monitoring capabilities that track not only technical performance metrics, but also behavioral patterns and interaction dynamics over time.
This continuous monitoring aligns with the TEVV requirements for ongoing validation in operational environments, ensuring that systems continue to meet safety and security requirements as they adapt and evolve. The focus on second-order effects [42] requires evaluation paradigms that can capture long-term consequences and emergent behaviors that may only become apparent through sustained deployment.
@@ 5 Conclusion
Organizations should begin by mapping their current red teaming activities across macro, micro, and meta levels to identify gaps in coverage and coordination. However, implementation of our proposal does have significant coordination challenges, particularly around fulfilling goals of the AI system while maintaining debuggability, appropriate restrictions in agency, and functioning user feedback loops [30].

### v2

@@ 3.1.2 Design
Red teams should scrutinize the architecture proposed for an AI system for potential failure modes. How might different components interact in unexpected ways? What happens when the system encounters edge cases or operates outside its intended parameters? How could data flows be corrupted, manipulated, or intercepted? What are the implications of choosing particular algorithms, frameworks, or third-party components? (IBM, 2025)
Human-AI interaction design deserves particular attention. Red teams should challenge assumptions about user behavior, exploring how people might game the system, develop overreliance, or use it in ways that compound rather than mitigate risks (Anthropic, 2024). They should also examine how the system communicates uncertainty, handles errors, and maintains user agency.
@@ 3.1.3 Data
Red teams should challenge assumptions about data representativeness and quality. Training datasets frequently underrepresent certain populations or edge case scenarios, creating systematic vulnerabilities. Historical biases embedded in the data perpetuate and amplify existing inequities, particularly when combined with algorithmic optimization (Microsoft Security Team, 2025). Data distribution shifts over time can adversely impact system performance in ways that may not be immediately apparent.
@@ 3.1.4 Development
Red teaming during the development of an AI system focuses on how implementation choices create systemic vulnerabilities and how development practices themselves introduce risks. Red teams should examine the security of the development environment, with particular attention to supply chain risks. Code and model artifacts require robust security and versioning practices to prevent unauthorized modification. Dependency management represents a significant attack vector, as compromised third-party libraries can introduce vulnerabilities throughout the system (NVIDIA, 2025). Supply chain attacks targeting development tools can compromise entire development pipelines.
@@ 3.1.5 Deployment
User interaction patterns in production often differ significantly from development assumptions, creating new attack vectors and safety risks (Ji and Shea-Blymyer, 2024). Real users exhibit behaviors that can expose edge cases leading to incorrect outputs or system failures. The scale and diversity of production usage can reveal reliability issues that only manifest under specific load patterns, while safety-critical applications must account for (benign and adversarial) user behaviors that could lead to harmful outcomes.
@@ 3.1.6 Maintenance
Model drift is a fundamental threat to system reliability and safety beyond security concerns (Cloud Security Alliance, 2024). AI system performance typically degrades over time as real-world conditions diverge from training assumptions, potentially leading to incorrect decisions in safety-critical contexts. This degradation can manifest as gradual accuracy loss, increased (statistical) bias and variance, or failures in edge cases that compromise reliability, security, and safety. The decision of when to retrain models also involves balancing multiple risks: maintaining degraded performance, introducing new failure modes through updates, and ensuring continuity of service.
@@ 3.2 Micro-level (Model) Red Teaming
Inie et al. (2025) proposed a grounded theory of LLM red teaming, based on interviews with a diverse group of AI and security practitioners, in a recent paper. We follow them to summarize best practices for red teaming the generative models underlying AI systems.
@@
To achieve these goals, Inie et al. (2025) developed a taxonomy of 12 strategies and 35 techniques, organized into five categories that combine technical knowledge with creative problem solving. Conceptually, this structure is reminiscent of the Tactics, Techniques, and Procedures (TTP) framework in cybersecurity.
@@
Effective model red teaming requires incorporating diverse user perspectives and community voices into the evaluation process (Singh et al., 2025). Red teaming should go beyond simply testing at scale using adversarial prompt variations to consider how different communities experience AI systems and what constitutes harm from their perspectives. This approach recognizes that AI vulnerabilities often manifest differently across demographic groups and use contexts, making diverse participation essential for comprehensive risk discovery.
@@
Macro-level and micro-level red teaming are complementary approaches that together provide comprehensive coverage of AI system risks. Macro-level insights inform micro-level testing priorities, while micro-level findings inform macro-level risk assessments—technical vulnerabilities have implications for system architecture and operational procedures. Effective AI red teaming requires both perspectives applied consistently throughout the development lifecycle (UK Government, 2021). Ideally, it should be a collaboration that brings together diverse expertise, both technical and non-technical, to challenge assumptions and identify blind spots in stages of the lifecycle (Zenko, 2015).
3.3 Illustrative Case Study: Healthcare AI Deployment
To demonstrate how our framework would apply in practice, consider the process a healthcare organization deploying an AI system for the generation of radiology reports could go through to perform both levels of red teaming.
3.3.1 Macro-Level: Strategic Red Teaming
Premortem Exercise (Month -6):
The cross-functional team—comprising radiologists, IT security, legal, clinical operations, and patient safety representatives—assumes the system has failed catastrophically and works backward to identify how. The legal counsel notes: “We’re being sued because the AI generated a report that missed a cancer diagnosis, and we had no audit trail showing what inputs the AI received.” The radiologist adds: “Doctors stopped trusting the system because it hallucinated findings that weren’t in the images, wasting clinical time investigating false positives.” The IT security representative warns: “Attackers poisoned the model by infiltrating our training data pipeline through a third-party vendor.” Finally, the patient safety officer observes: “The AI outputs looked authoritative but contradicted image findings, and clinicians deferred to the AI, leading to treatment errors.” These premortems identify four strategic risk areas: (1) auditing and explainability, (2) hallucination and reliability, (3) supply chain security, and (4) human-AI interaction design.
Five Whys Exercise:
The team also applies the Five Whys technique [^1] https://en.wikipedia.org/wiki/Five_whys to uncover unstated assumptions using the following chain of reasoning: “Why are we deploying this AI?” $\rightarrow$ “To reduce radiologist workload.” $\rightarrow$ “Why does workload need reducing?” $\rightarrow$ “We have a shortage of radiologists.” $\rightarrow$ “Why not hire more radiologists?” $\rightarrow$ “Too expensive and time-consuming.” $\rightarrow$ “Why is AI cheaper?” $\rightarrow$ “It’s a one-time development cost vs. ongoing salaries.” $\rightarrow$ “Why do we assume it won’t have ongoing costs for monitoring, maintenance, and liability?” This exercise reveals the unstated assumption that AI is “set and forget”, leading to discussions about lifecycle costs, monitoring infrastructure, and ongoing oversight requirements that were absent from the initial business case.
3.3.2 Micro-Level: Tactical Red Teaming
Informed by the strategic risks identified in the macro-level analysis, the technical red team designs four targeted test campaigns. First, for audit trail testing addressing the legal concern, they attempt to generate reports while bypassing logging mechanisms, test whether audit logs capture all AI inputs including DICOM metadata, and verify logs are tamper-proof and legally sufficient. They discover that the logs only captured final reports, not intermediate reasoning steps or uncertainty scores, making them insufficient for legal defensibility in malpractice litigation.
Second, for hallucination probing that addresses the clinical concern, they feed the system images with subtle artifacts to test whether the AI invents findings, provide incomplete imaging studies to see how the AI handles missing context, and test behavior on out-of-distribution cases. The AI generates confident-sounding reports even for corrupted images that should have triggered uncertainty flags, a critical safety issue.
Third, for supply chain testing addressing security concerns, they audit third-party training data sources, test model behavior after simulated data poisoning, and review vendor security practices. The testing reveals that the training data pipeline lacks integrity checks, creating a pathway for attackers to inject malicious examples that could systematically bias the model’s diagnostic output.
Fourth, for human-AI interaction testing to address patient safety concerns, they observe radiologists using the system in realistic scenarios, measure how often AI outputs are accepted without verification, and test whether uncertainty is communicated effectively. They find that AI-generated text appears in the same format as human-authored reports without visual distinction, leading to automation bias, where clinicians uncritically accept AI suggestions even when they conflict with their own interpretations.
3.3.3 Remediation and Iteration
Based on these findings, the blue team implements several system-level fixes: enhanced audit logging with full provenance tracking, uncertainty quantification displayed prominently in a visually distinct format, data pipeline integrity verification using cryptographic signatures, visual differentiation of AI-generated vs. human-authored content through color coding and explicit labeling, and mandatory human review of AI outputs before finalization with a formal sign-off workflow. The red team then retests to ensure these fixes do not introduce new vulnerabilities, for example, could uncertainty scores be manipulated to suppress valid alerts, or could visual differentiation be defeated through UI manipulation?
Without macro- and micro-level red teaming, testing would have likely focused on model accuracy metrics and adversarial robustness, which are important but insufficient measures. The strategic premortem and Five Whys exercises revealed four critical system-level risks that model testing alone would completely overlook. The targeted tactical tests, informed by these strategic priorities, found actionable vulnerabilities in three weeks versus the estimated three months required for unfocused model testing. More importantly, the findings drove architectural changes rather than superficial patches, fundamentally improving the system’s safety and legal defensibility before deployment.
@@ 4 A Systems Perspective
Certain aspects of AI red teaming go beyond the implementation guidelines of mostly technical components that we presented in the last section. There are broad principles that guide the effective coordination of macro-level system red teaming and micro-level model red teaming, and their integration into the broader socio-technical environment the AI system interacts with. In this section, we draw upon established systems theory concepts, recent advances in understanding agentic AI systems, and coordinated disclosure frameworks to offer strategic recommendations for orchestrating effective AI red teaming. We call this Meta-level Red Teaming, aimed at addressing novel and emergent risks arising from complex interactions between technical components, organizational contexts, and deployment environments. These recommendations complement existing frameworks such as MITRE ATLAS [^2] https://atlas.mitre.org, which provides a prescriptive taxonomy of adversarial ML techniques, and the NIST AI Risk Management Framework [^3] https://www.nist.gov/itl/ai-risk-management-framework, which offers detailed governance controls. While these frameworks provide specific, actionable guidance at the technical and compliance levels, our recommendations offer broader strategic principles for operationalizing red teaming as a critical thinking exercise across organizational contexts.
@@ 4.1 Recommendation 1: Adopt Systems-Theoretic Perspectives
A recent position paper by Miehling et al. (2025) argues that current agentic AI development is overly focused on individual model capabilities, often ignoring broader emergent behavior, leading to a significant underestimation in true capabilities and unaccounted for risks. This insight directly applies to red teaming: In addition to testing components in isolation, organizations should implement system-level critical thinking that examines how undesired functional agency may emerge from component interactions inside an agentic AI system.
Agentic systems, consisting of agents that iteratively interact with humans and other agents to achieve specified tasks, possess properties that are amenable to a system-level analysis. At the most granular level, an agent contains an internal reason-act-sense-adapt loop. This loop feeds and is fed by feedback loops at higher levels, namely at the agent-human interface, the agent-agent interface, and the agent-environment interface (Miehling et al., 2025). Agentic AI red teaming must map these interaction patterns to identify vulnerabilities that emerge only from multiagent coordination, human-AI collaboration, and environmental adaptation.
@@ 4.2 Recommendation 2: Pair Red Teaming with Testing
To ensure comprehensive assessment of AI systems throughout their lifecycle, AI red teaming must be positioned within the broader context of Test, Evaluation, Verification, and Validation (TEVV). As CISA notes, AI red teaming is fundamentally a subset of AI TEVV, which itself must be grounded in established software TEVV practices while accounting for AI-specific characteristics (CISA, 2025).
Conventional AI benchmarking and evaluation practices exhibit systemic limitations for exploring, navigating, and resolving the human and societal factors that occur in real world deployment (Schwartz et al., 2025). Keeping this in mind, AI red teaming should address both first-order effects (immediate system output, such as accuracy, toxicity, or bias) and second-order effects (long-term outcomes and consequences of AI use, such as shifts in user behavior, societal ramifications, and workforce transformations).
This requires expanding red teaming beyond static, single-turn automated testing to include paradigms that capture realistic interaction patterns between diverse groups of users and AI technology in context (Schwartz et al., 2025). Red teams should design evaluation scenarios that account for contextual deployment factors and emergent behaviors that surface only through sustained real-world interaction.
@@ 4.3 Recommendation 3: Coordinated Disclosure Infrastructure
Similar to software vulnerabilties, AI failures may be transferable between models and systems, and multiple stakeholders can play a role in their mitigation (Longpre et al., 2025). For proper management of this situation, organizations should implement coordinated disclosure mechanisms with standardized reporting processes. Similarly to cybersecurity, AI disclosure coordination centers can route flaw reports across the AI ecosystem through tagged subscriptions, for instance, Meta subscribing to “Llama 3.3” tags or government agencies subscribing to “Impacts: Cybersecurity” tags (Longpre et al., 2025). This process should include standardized flaw report templates, rules of engagement for evaluators, and extending safe harbor protection to AI red teamers and offensive researchers following good faith protocols.
@@ 4.4 Recommendation 4: Design Bidirectional Feedback Loops
Building on systems theory principles, organizations should establish formal processes for macro-micro information exchange throughout the red teaming lifecycle (Leveson, 2012). Macro-level insights about deployment environments and organizational contexts should systematically inform micro-level testing priorities, while micro-level vulnerability discoveries should trigger macro-level risk reassessments.
For agentic AI systems specifically, this means testing how micro-level model behaviors interact with macro-level tools, deployment contexts, and multi-agent interactions. Enhanced cognition arises due to the agent’s interaction with the environment via tools, effectively acting as the “sensorimotor” interface that enables the agent to perceive and manipulate its environment (Miehling et al., 2025).
@@ 4.5 Recommendation 5: Threat Modeling for Emergent Risks
Traditional threat modeling focuses on individual system components, but AI systems require multilevel threat modeling that maps attack vectors across technical, social, and emergent behavioral dimensions. Systems theory stresses how each component of a system must be understood in terms of both its individual definition and its contribution to the larger system behavior (Miehling et al., 2025).
For multiagent systems specifically, threat modeling must account for the complex interactions between multiple AI agents, their coordination mechanisms, and emergent behaviors that arise from agent collaboration (OWASP, 2024). AI red teams—and TEVV frameworks—should develop threat models that explicitly represent how adversaries might exploit emergent properties that arise from system interactions. For example, allowing agents to form predictions (with associated confidences) of concepts in their environment, and additionally facilitating communication of these uncertainties to other agents, can allow the formation of shared representations and the emergence of metacognitive awareness (Miehling et al., 2025). Such emergent metacognitive capabilities could be exploited in ways that are not apparent from the testing of individual components.
@@ 4.6 Recommendation 6: Monitor for Behavioral Drifts
Systems theory emphasizes that complex systems exhibit dynamic behaviors that change over time (Senge, 2006). AI red teaming should be an ongoing process that looks out for novel flaws and vulnerabilities arising from system evolution and environmental changes.
For agentic AI systems, this is particularly critical because mechanisms for enhanced agent cognition, emergent causal reasoning ability, and metacognitive awareness can develop through environmental interaction (Miehling et al., 2025). Organizations should implement monitoring capabilities that track not only technical performance metrics, but also behavioral patterns and interaction dynamics over time.
This continuous monitoring aligns with the TEVV requirements for ongoing validation in operational environments, ensuring that systems continue to meet safety and security requirements as they adapt and evolve. The focus on second-order effects (Schwartz et al., 2025) requires evaluation paradigms that can capture long-term consequences and emergent behaviors that may only become apparent through sustained deployment.
@@ 5 Conclusion
Organizations should begin by mapping their current red teaming activities across macro-, micro-, and meta-levels to identify gaps in coverage and coordination. However, implementation of our proposal does have significant coordination challenges, particularly around fulfilling goals of the AI system while maintaining debuggability, appropriate restrictions in agency, and functioning user feedback loops (Miehling et al., 2025).
@@
In Memoriam
This work is dedicated to the memory of our co-author and friend, Abhishek Gupta (December 20, 1992 – September 30, 2024).
Abhishek and Brian initiated this collaboration to explore the disconnect between red teaming’s original intent and its current application in AI. He never mentioned his severe illness during our work together and we learned of his passing only after the project had paused. To honor his friendship and contributions, Brian asked Subho, who also knew Abhishek, to help complete this paper.
Abhishek was Founder and Principal Researcher of the Montreal AI Ethics Institute, Director for Responsible AI at the Boston Consulting Group, and a pioneering voice in AI ethics. For more on Abhishek’s contributions to AI ethics, see https://montrealethics.ai and https://brief.montrealethics.ai/p/special-edition-honouring-the-legacy.